OpenStudy (anonymous):

what technique would be best applied to factor the polynomial below. x^2-25

OpenStudy (anonymous):

(a^2-b^2)

OpenStudy (anonymous):

here a=x and b=25

OpenStudy (anonymous):

(a^2-b^2)=a^2-2ab+b^2 use this

OpenStudy (cwrw238):

this is a form of quadratic and is the product of 2 binomials also its difference of 2 squares as qib as indicated plug in values of a and b and you've got it

OpenStudy (anonymous):

so you have to group them?

OpenStudy (cwrw238):

actually b = 5 not 25

OpenStudy (cwrw238):

a^2 - b^2 = (a - b)( a + b)

OpenStudy (anonymous):

yeah sorry b=5

OpenStudy (anonymous):

im sorry this is the right formula a^2 - b^2 = (a - b)( a + b)

OpenStudy (anonymous):

so its grouping?

OpenStudy (anonymous):

a^2 - b^2 = (a - b)( a + b) (x)^2-(5)^2=(x-5)(x+5)

OpenStudy (anonymous):

ok